OpenStudy (anonymous):

what is the inverse of f(x)=14.95 + 3x

OpenStudy (sleepyhead314):

Hello and Welcome to OpenStudy! :) to find the inverse of a function, switch x and y so y = 14.95 + 3x becomes x = 14.95 + 3y now solve for y from there, do you know how to do that?

OpenStudy (anonymous):

i did that part already, I'm stuck at what to do after that

OpenStudy (sleepyhead314):

ok, so we want y all alone on one side so first subtract 14.95 from both sides then divide both sides by 3

OpenStudy (anonymous):

ok so, I'm supposed to have x-14.95/3 =y??

OpenStudy (sleepyhead314):

mmm either put a parenthesis around it like (x-14.95)/3 or distribute it so (1/3)x - (14.95)/3 = y

OpenStudy (anonymous):

so thats the answer for the inverse? no more steps?

OpenStudy (sleepyhead314):

yep :)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

wait i have an addition problem related to that

OpenStudy (anonymous):

find f ^-1 (38.95)

OpenStudy (sleepyhead314):

so once you got f^-1(x) = (x-14.95)/3 plug in 38.95 for x

OpenStudy (anonymous):

ok how do you find the range of the inverse function of f(x)=ln(x-5)+2

OpenStudy (sleepyhead314):

annnnnddddd that is something that I've deleted from my memory, sorry ummm try this http://www.wolframalpha.com/input/?i=f%28x%29%3Dln%28x-5%29%2B2+range but I won't be able to explain it for you, sorry

OpenStudy (anonymous):

how bout the y intercept of f(x)=ln(x-5)+2 , its fine if you don't know

OpenStudy (sleepyhead314):

ahhhh mk that's something I kinda know the parent function is y = ln(x) how it is moved around is y = ln(x-h) + k where h is how much left <-> right and k is how much up down orrrr you can think y = mx + b and b = y-intercept same idea for this format... so 2 is the y-intercept

OpenStudy (anonymous):

i have to show my work , so can you explain a little more how you arrived at 2

OpenStudy (sleepyhead314):

oh... show work... thennnn y-intercept is when x = 0 plug in 0 for x and simplify

OpenStudy (anonymous):

k thanks

OpenStudy (sleepyhead314):

*scratches head* I guess it's not 2 then... *shrugs* I failed algebra 2 xD but the thing about y-intercept is when x=0 should be right